OpenStudy (anonymous):

Choose the point slope form of the question below that represents the line that passes through the point (-1,6) and has a slope of -3 A: y-6=-3x-3 B: y-6=-3(x+1) C: y=-3x+3 D: 3x+y=3

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

okay do you know what point slope form is?

OpenStudy (anonymous):

@atreyu6s3x6 is that y-y or something like that omg i forgot

OpenStudy (anonymous):

yea you got the right idea, its y – y1 = m(x – x1)

OpenStudy (anonymous):

@atreyu6s3x6 aah thank you

OpenStudy (anonymous):

yea so do you know what to do now?

OpenStudy (anonymous):

@atreyu6s3x6 plug in?

OpenStudy (anonymous):

yea.. so try it and let me know what you get

OpenStudy (anonymous):

okay, y-6 =-3(x-1) ?

OpenStudy (anonymous):

close but dont forget that a negative and a negative are a positive, so -1 changes to a positive 1

OpenStudy (anonymous):

@atreyu6s3x6 oh thats right whoops

OpenStudy (anonymous):

yea but nice try